Calculate the pH and the concentrations of all species present (H3O+, F-, HF and OH-) in 0.05 M HF.
Answer #1

Let α be the dissociation of the weak acid
                            HF + H2O <---> H3O + + F-

initial conc.            c 0 0

change                -cα +cα +cα

Equb. conc.         c(1-α) cα cα

Dissociation constant , Ka = cα x cα / ( c(1-α)

                                         = c α2 / (1-α)

In the case of weak acids α is very small so 1-α is taken as 1

So Ka = cα2

==> α = √ ( Ka / c )

Given Ka = 6.6x10-4

          c = concentration = 0.05 M

Plug the values we get α = 0.1149

[H3O+] = cα = 0.05x0.1149 = 5.74x10-3 M

[ F-] =  cα = 0.05x0.1149 = 5.74x10-3 M

pH = - log [H3O+]

    = - log( 5.74x10-3)

    = 2.24

[HF] =  c(1-α) = 0.044 M

[OH-] = (1.0x10-14) /  [H3O+] = 1.74x10-12 M